### v2

@@
(April 1, 2025)
This paper presents the first quasi-experimental research examining the effect of both local and state anti-discrimination laws on sexual orientation on the labor supply and wages of lesbian, gay, and bisexual (LGB) workers. To do so, we use the American Community Survey data on household composition to infer sexual orientation and combine this with a unique panel dataset on state and local anti-discrimination laws. Leveraging variation in law implementation across localities over time and between same-sex and different-sex couples, we find that anti-discrimination laws not only significantly narrow gaps in labor force participation and employment for male same-sex couples relative to men in different-sex couples, but also boost their percentile rank in the wage distribution. Our analysis reveals mostly null effects for female same-sex couples; however, in metropolitan areas these laws significantly reduce their employment compared to women in different-sex couples. One explanation for the reduced labor supply is that female same-sex couples begin to have more children in response to the laws. Finally, we present evidence that state anti-discrimination laws significantly and persistently increased support for same-sex marriage. This research shows that anti-discrimination laws can be an effective policy tool for reducing labor market inequalities across sexual orientation and improving sentiment toward LGB Americans.
@@ 1 Introduction
On April 15, 2020, the Supreme Court of the United States ruled that federal employment discrimination law applied to lesbian, gay, bisexual, and transgender (LGBT) individuals. Before this ruling, the legislation in 25 states made it legal for private businesses to fire an employee for being LGBT. Unlike race, sex, age, and religion, federal law did not include sexual orientation as a protected class for federal anti-discrimination laws until 2020. As a result, laws extending protected class status based on sexual orientation varied between states and localities. The first state to pass an anti-discrimination law based on sexual orientation was Wisconsin in 1982, and the most recent state to pass such a law was Pennsylvania and Michigan in 2023.111Since August 2018, Pennsylvania has interpreted discrimination based on sexual orientation and gender identity as being covered under the “sex” category of the Pennsylvania Human Relations Act. This allowed LGBT individuals to file complaints for discrimination in employment, housing, education, and public accommodations. Following Michigan’s lead in May 2018, Pennsylvania became the second state to implement statewide LGBT protections in this manner. There has been a renewed push to enact sexual orientation protections at the legislative level as Democrats in the House of Representatives introduced “The Equality Act” in early 2019, which would add sexual orientation and gender identity as federally protected characteristics. The focus of this paper is the effect of local and state sexual orientation anti-discrimination legislation on labor market differentials for lesbian, gay, and bisexual (LGB) workers. [^2] For a review of the impacts of legal access to same-sex marriage, see Badgett et al. (2024). For a broader review of the economics related to sexual orientation and gender identity, see Badgett et al. (2023).
@@
We exploit the differential roll-out of state and local laws from 2005-2019 in a modern triple difference-in-differences framework to analyze how these anti-discrimination laws impact the wages and labor supply of LGB workers. The economics literature on LGB workers consistently finds that gay/bisexual men have a pay and labor supply penalty and lesbian/bisexual women have a pay and labor supply premium over their heterosexual counterparts333(Badgett (1995); Klawitter (2011); Klawitter and Flatt (1998); Allegretto and Arthur (2001); Black et al. (2003); Carpenter (2005); Black et al. (2007); Antecol et al. (2008); Klawitter (2015); Jepsen and Jepsen (2017); Carpenter and Eppink (2017)), Tilcsik et al. (2015). The differences in pay are attributed to a host of factors including discrimination and intra-household labor allocation and specialization. We test the role of discrimination by examining how the passage of anti-discrimination laws affects the labor supply and wage differences between LGB and heterosexual workers.
There have been previous studies examining how these state and local laws affect the pay gap, but these studies have been limited in their ability to identify causal effects (Klawitter and Flatt (1998); Gates (2009)). Klawitter and Flatt (1998) compared people in same-sex partnerships to those in different-sex partnerships in the 1990 Census and found that same-sex couples have no significant difference in wages in places with anti-discrimination laws. Gates (2009) uses the 2000 Census and conducts a similar analysis. Gates (2009) finds that gay/bisexual men in places with anti-discrimination laws have a 3% wage premium over gay/bisexual men in places without these laws, and lesbian/bisexual women have a 2% wage premium over lesbian/bisexual women in places without these laws [^4] Tilcsik (2011) found that resumes signaling LGB status received significantly fewer callbacks in localities without sexual orientation anti-discrimination laws. These past studies on anti-discrimination laws, while informative, fail to fully account for changes over time.
Previous research has leveraged state-level sexual orientation anti-discrimination laws within a difference-in-differences framework. [^5] The Williams Institute found that sexual orientation discrimination occurs at a similar rate to sex-based discrimination using state-level complaint data following state anti-discrimination laws (Ramos et al., 2008). Martell (2013) study demonstrated that such laws have significantly reduced wage disparities for gay men by 20%.666 Since Martell (2013) and Klawitter and Flatt (1998) there has been a booming literature on difference-in-differences. Accordingly, unlike Martell (2013) and Klawitter and Flatt (1998), we employ modern difference-in-difference estimators to alleviate biases stemming from staggered policy design. Our research extends beyond the scope of previous studies by incorporating an examination of local laws, thereby offering a more comprehensive understanding of the legislative landscape. Furthermore, this study also estimates the effects of such laws on women, a demographic previously under-explored in this context. Our study expands the scope of Klawitter and Flatt (1998) by leveraging a larger dataset, utilizing American Community Survey (ACS) data spanning from 2005 to 2019, in contrast to Klawitter and Flatt (1998) which relied on the 1990 U.S. census data encompassing approximately 13,000 married couples, 14,500 different-sex unmarried couples, and 6,800 same-sex couples. The ACS data allows for a robust analysis across a broader temporal spectrum, capturing a wide array of demographic and economic changes over 15 years. This extensive dataset, combined with the application of modern difference-in-difference estimators, enables a more detailed and dynamic examination of the effects of anti-discrimination policies on the earnings and household incomes of same-sex couples. Additionally, Burn (2018)’s more recent examination in a triple difference-in-differences framework confirms these findings but also underscores a gap in the literature: the lack of consideration for local anti-discrimination ordinances. Previous studies have not fully addressed potential biases arising from variations in the timing of these laws’ enactment.
This paper is the first to examine both local and state-level sexual orientation anti-discrimination laws in a quasi-experimental design. This paper is also the first to analyze the effects of any local anti-discrimination laws on sex, race, or sexual orientation in a modern difference-in-difference framework [^7] Using state variation in anti-discrimination laws has also been analyzed to understand racial- and sex-based discrimination (Neumark and Stock (2006); (Donohue III and Heckman, 1991);((Margo, 1995); Goldin and Margo (1992)). However, these studies on sex and race anti-discrimination laws also omit local laws, potentially skewing their estimates..
@@
We use the 2005-2019 American Community Survey (ACS) and household composition to infer sexual orientation and create a unique and novel panel dataset on the passage of local and state anti-discrimination laws. We collected information on local laws from a host of sources including media reports, FOIA requests, and an advocacy group. We find a significant reduction in differences between LGB workers and heterosexual workers across labor supply and wage measures, due to anti-discrimination laws. Anti-discrimination laws significantly reduce the gap in labor force participation and employment of men in male same-sex couples by 3.8 percentage points (78%) and 3.7 percentage points (77%), respectively. The laws improve men in male same-sex couples relative position in the wage distribution, with their percentile rank in the hourly wage and annual earnings distributions improving on average by approximately 2.7 and 3 percentiles, respectively. By contrast, the wage distribution for women in same-sex partnerships remains largely unaffected by these laws. Using an event study plot, we show that the outcome trends are parallel before the implementation of the anti-discrimination laws suggesting the workers in treatment and control regions are reasonable comparisons. [^8] In Appendix B A, we replicate the findings of past research that finds men in male same-sex couples have lower labor force participation and employment rates and make 8-11% less than their employed straight counterparts using hourly wages and annual earnings. We also replicate findings that women in female same-sex couples have a higher labor force participation and employment rate and earn 5-15% more than their employed straight counterparts.
@@
We explore theories for the differing effects of anti-discrimination laws on men in male same-sex couples and women in female same-sex couples in the Discussion section. Reduced discrimination improves job security for sexual minorities (Badgett, 2006), which may allow the secondary earner to reduce market participation in favor of increased household production, including child rearing, a pattern consistent with household bargaining models (Chiappori, 1992; Lundberg and Pollak, 1993). These differences in household specialization are less pronounced in same-sex partnerships, but they may become more similar for women in female same-sex couples following the passage of an anti-discrimination law if it gives greater protection to the higher wage earner. We show empirically that female same-sex households have significantly more children than male same-sex households after the passage of an anti-discrimination law.
@@
Finally, our last contribution is the use of polling data on the support for same-sex marriage to examine the relationship between anti-discrimination laws and state sentiment toward LGB workers. One may expect that sentiment toward LGB workers would increase right before the passage of the laws, creating a selection issue. Alternatively, the anti-discrimination laws may normalize being a sexual minority, and improve sentiments toward LGB workers. We collected every poll on same-sex marriage for each state and year from the Pew Research Center. We show that the passage of state-wide sexual orientation anti-discrimination laws persistently increases favorability toward LGB people through increased support for same-sex marriage. The states that passed anti-discrimination laws had parallel pre-trends in support of same-sex marriage before the passage of the laws. Afterward, those states had a significant and persistent increase in their support for same-sex marriage. The observed increase in support following the enactment of the law, rather than before, implies that policy changes can indeed influence public opinion. This notion aligns with the findings of Sansone (2019), suggesting that legislative actions can lead to shifts in societal attitudes. This is in contrast to the thermostatic model proposed by Wlezien (1995), which posits that public sentiment drives policy changes rather than the other way around. Increased favorability toward LGB people may act in conjunction with greater job security to improve labor market outcomes for LGB Americans.
@@ 2 Data
A common issue in analyzing the pay disparities between homosexual and heterosexual workers is a lack of high-quality data about sexual orientation, wages, and employment. We follow the literature (Klawitter and Flatt (1998); Gates (2009); Jepsen and Jepsen (2017)) in inferring sexual orientation by looking at household composition. Specifically, we infer a person’s sexual orientation to be homosexual/bisexual if they have an “unmarried partner” or a spouse that is the same sex as themselves. [^9] We exclude individuals whose sex or relationship status was allocated following (Gates and Steinberger, 2009).. The ACS references a close personal relationship for unmarried partners as well as giving an option for “partner/roommate”, which reduces the likelihood that different-sex roommates would misclassify as unmarried partners.
The comparisons that we make are between people in same-sex partnerships (SSP), defined as being in an unmarried partnership or married with someone of the same sex, and those in different-sex partnerships (DSP). This comparison based on household composition, though standard in the literature, is not equivalent to comparing LGB and heterosexual workers. It compares those who are in a same-sex relationship to those who are in a different-sex relationship. Importantly, this comparison excludes all workers who are not in a cohabitating relationship and classifies bisexual individuals in different-sex relationships as members of the control group. Because bisexual workers consistently earn less than both heterosexual and gay or lesbian counterparts ((Mize, 2016)), this classification introduces a bias toward underestimating pay disparities. Additionally, similar misclassification may occur if gay or lesbian individuals are categorized as heterosexual. Thus, the estimated pay gaps presented here likely represent conservative estimates.
@@
We use the 2005-2019 yearly ACS from (Ruggles et al., 2019). We use wages, defined as salaried wages from an employer. The ACS reports wage earnings in annual terms. We use the annual figures and convert them to hourly wage figures using variables on average weekly hours worked and weeks worked101010“Weeks worked” in the ACS is a categorical variable giving a range of weeks worked. We take the median value given in the range to compute weeks worked.. Finally, we limit our sample to prime-age working adults and only examine those who are ages 25-65. We use 25 as the lower age cutoff to allow workers to complete college and have more fully developed human capital, which is typically thought of as an important factor for the differences in wages between LGB and straight workers (Black et al., 2007).
@@
We obtained data on the passage of state anti-discrimination laws from LGBTMap.org, an LGBT advocacy group. The website gives information on which states passed anti-discrimination laws and when. We focus solely on sexual orientation anti-discrimination laws that give protection in employment. LGBTMap.org also provided incomplete data detailing the passage of local city and county laws with many cities missing years for the start of their anti-discrimination laws. We completed the dataset for the relevant years using old media reports, correspondence with local officials, and FOIA requests. Building on earlier city-level analyses such as (Klawitter, 2011), this dataset provides the first complete city-level dataset covering 15 years of sexual orientation anti-discrimination laws in the U.S. We focus on cities reported in the ACS and match those city laws with their corresponding counties to merge with the ACS. The ACS only reports county of residence for those in metro areas, so any rural counties are lumped together within a state.
We report the state, city, and county laws and the year they were enacted in Table 1. In Figure 1, we show which counties had sexual orientation anti-discrimination laws in 2005 and 2019, respectively. In our analysis, laws are coded based on their initial application to private employment and housing, focusing on broad protections rather than sector-specific ones. For instance, while Delaware extended protections to state employees in 1999, our coding reflects 2009, the year broader protections were enacted. This approach could overlook the incremental impact of earlier, narrower protections on societal attitudes. The same principle applies to local laws, where initial protections for public employees in some cities might precede more comprehensive legislation, potentially influencing early shifts in public sentiment towards sexual minorities. The enforcement mechanisms of anti-discrimination laws, such as fines, mandatory training, and public rebukes, serve as significant deterrents, encouraging the adoption of more inclusive employment policies. The rise in successful legal challenges against companies concerning discrimination highlights the real legal and financial consequences they encounter, reinforcing the laws’ effectiveness in promoting workplace equality (for Human Resource Management (2023)).
The enactment of anti-discrimination legislation is not random. The states that have these protections are generally considered friendlier to LGB workers than those without these laws and have a larger share of same-sex partnerships than those states without these laws. Also, many counties with anti-discrimination laws have large cities with a larger concentration of LGB workers than rural counties. However, this is not universal. For instance, Utah extended protection to LGB workers despite being a relatively conservative state, and certain liberal cities like Houston [^11] Houston passed a sexual orientation and gender identity anti-discrimination ordinance, but it was only in effect for 3 months before being challenged. The law was put up to a public vote and lost, repealing the law. are noticeably absent from extending protection to LGB workers. One potential concern in this analysis is the endogenous adoption of laws. The areas that are friendly and less hostile to LGB workers may be the areas that are most likely to adopt sexual orientation anti-discrimination laws. The main threat to identification is that the timing of anti-discrimination laws is correlated with another factor causing an omitted variable bias. It could be that positive sentiment toward LGB workers causes both the law change and any change in labor market outcomes. We mitigate this by adopting a saturated triple-difference model, including all two-way fixed effects (county-year, year-SSP, and county-SSP), as proposed by (Burn, 2018). The inclusion of county-year fixed effects remove the need for county controls, addressing the lack of local support data for same-sex marriage. We refine the approach of (Burn, 2018) by employing the methodology of (Borusyak et al., 2021), designed for analyzing staggered policy adoptions like ours. This method accurately estimates the effects of anti-discrimination laws, considering their staggered implementation times.
In Figure 2, we present the distribution of same-sex partnerships by state from 2005 to 2019, and in Table 2, we present the 10 counties with the largest share of same-sex partnerships. Unsurprisingly, the distribution of same-sex partnerships by state and by county is skewed toward more progressive states and counties with large cities in them that are known for having a large LGB population like San Francisco, the District of Columbia, New York City, and Boston. The LGB population varies from the heterosexual population in many dimensions including geography.

### v3

@@
(November 19, 2025)
This paper presents quasi-experimental research examining the effect of both local and state anti-discrimination laws on sexual orientation on the labor supply and wages of lesbian, gay, and bisexual (LGB) workers. To do so, we use the American Community Survey data on household composition to infer sexual orientation and combine this with a unique panel dataset on state and local anti-discrimination laws. Leveraging variation in law implementation across localities over time and between same-sex and different-sex couples, we find that anti-discrimination laws significantly narrow gaps in labor force participation and employment for men in same-sex couples relative to men in different-sex couples, and also increase their percentile rank in the wage distribution. Our analysis reveals mostly null effects for female same-sex couples; however, in metropolitan areas these laws significantly reduce their employment compared to women in different-sex couples. One explanation for the reduced labor supply is that female same-sex couples begin to have more children in response to the laws. Finally, we present evidence that state anti-discrimination laws significantly and persistently increased support for same-sex marriage. This research shows that anti-discrimination laws can be an effective policy tool for reducing labor market inequalities across sexual orientation and improving sentiment toward LGB Americans.
@@ 1 Introduction
On June 15, 2020, the U.S. Supreme Court held in Bostock v. Clayton County, 590 U.S. 644 (2020), that Title VII’s prohibition of sex discrimination in employment includes discrimination based on sexual orientation and gender identity. Prior to this ruling, more than half of U.S. states lacked explicit statewide employment protections based on sexual orientation, though many localities had adopted their own ordinances. As a result, protections varied substantially across and within states throughout the period studied. Wisconsin became the first state to enact a statewide sexual orientation employment non-discrimination law in 1982. There has been continued legislative interest in expanding federal protections, most notably the introduction of the “Equality Act” in early 2019, which would add sexual orientation and gender identity to federal civil rights statutes.
This variation in state and local protections provides the basis for our empirical analysis. In this paper, we examine how sexual orientation employment non-discrimination laws affect labor market differentials for lesbian and gay workers.
@@
We exploit the differential roll-out of state and local laws from 2005–2019 in a modern difference-in-differences-in-differences (DDD) framework to analyze how these anti-discrimination laws impact the wages and labor supply of LGB workers. The economics literature consistently finds that gay men (or men in male same-sex couples) face earnings and labor-supply penalties, while lesbian women (or women in female same-sex couples) exhibit earnings and labor-supply premia relative to comparable heterosexual counterparts. [^1] For a review of the impacts of legal access to same-sex marriage, see badgett2024review. For a broader review of the economics related to sexual orientation and gender identity, see badgett2023review. In addition, see badgett1995wage; klawitter1998effects; allegretto2001empirical; black2003earnings; carpenter2005self; antecol2008sexual; klawitter2011multilevel; klawitter2015meta; jepsen2017self; carpenter2017does; Burn2018. The differences in pay are attributed to a host of factors including discrimination and intra-household labor allocation and specialization. [^2] Occupational sorting and concealment may also mediate gaps, with sexual minorities selecting into jobs with lower customer discrimination or greater autonomy (tilcsik2015concealable; black2007economics). Consistent with administrative complaints, sexual-orientation discrimination occurs at rates comparable to sex-based discrimination (ramos2008evidence). We test the role of discrimination by examining how the passage of anti-discrimination laws affects the labor supply and wage differences between LGB and heterosexual workers.
This paper provides quasi-experimental evidence on both local and state sexual orientation anti-discrimination laws. It is also, to our knowledge, the first study to evaluate the effects of local anti-discrimination ordinances, whether based on sex, race, or sexual orientation, within a modern difference-in-differences framework. [^3] Legal changes and enforcement have been used to study race- and sex-based discrimination (e.g., neumark2006labor; see also donohue1991continuous). These analyses do not incorporate local ordinances, which can bias estimates that rely on state-only policy variation.
@@
We use the 2005-2019 American Community Survey (ACS) and household composition to infer sexual orientation and create a unique and novel panel dataset on the passage of local and state anti-discrimination laws. We collected information on local laws from a host of sources including media reports, FOIA requests, and an advocacy group. We find a significant reduction in differences between LGB workers and heterosexual workers across labor supply and wage measures, due to anti-discrimination laws. Anti-discrimination laws reduce the labor force participation and employment gaps for men in male same-sex couples by 3.8 percentage points (78% of the baseline gap) and 3.7 percentage points (77%), respectively. The laws improve men in male same-sex couples’ relative position in the wage distribution: their percentile ranks in the hourly wage and annual earnings distributions rise by about 2.6 and 2.9 percentile points, respectively. By contrast, the wage distribution for women in same-sex couples remains largely unaffected by these laws. Event-study estimates show no pre-trends, supporting the parallel-trends assumption and the comparability of treated and control regions.
@@
We explore potential mechanisms for the differing effects on men and women in the Discussion section, without imposing a single unified explanation; as in the broader literature, some differences remain an empirical puzzle. Reduced discrimination improves job security for sexual minorities (badgett2006discrimination), which may allow the secondary earner to reduce market participation in favor of increased household production, including child rearing, a pattern consistent with household bargaining models (chiappori1992collective; lundberg1993separate). These differences in household specialization are less pronounced in same-sex couples, but they may become more similar for women in female same-sex couples following the passage of an anti-discrimination law if it gives greater protection to the higher wage earner. We show empirically that female same-sex households have significantly more children than male same-sex households after the passage of an anti-discrimination law.
@@
Finally, our last contribution is the use of polling data on the support for same-sex marriage to examine the relationship between anti-discrimination laws and state sentiment toward LGB workers. One may expect that sentiment toward LGB workers would increase right before the passage of the laws, creating a selection issue. Alternatively, the anti-discrimination laws may normalize being a sexual minority, and improve sentiments toward LGB workers. We compiled state–year polls on same-sex marriage from the Pew Research Center. We show that the passage of statewide sexual orientation anti-discrimination laws persistently increases favorability toward LGB people through increased support for same-sex marriage. The states that passed anti-discrimination laws had parallel pre-trends in support of same-sex marriage before the passage of the laws. Afterward, those states had a significant and persistent increase in their support for same-sex marriage. The observed increase in support following the enactment of the law, rather than before, implies that policy changes can indeed influence public opinion. This matches the findings of sansone2019pink, suggesting that legislative actions can lead to shifts in societal attitudes. This is in contrast to the thermostatic model proposed by wlezien1995public, which posits that public sentiment drives policy changes rather than the other way around. Increased favorability toward LGB people may act in conjunction with greater job security to improve labor market outcomes for LGB Americans.
Our study complements and extends prior state-focused analyses (e.g., klawitter2011multilevel; martell2013endas; Burn2018) by introducing a novel city–county ENDA panel and applying modern staggered-adoption estimators. Unlike work centered on marriage laws (e.g., hansen2020labor; hansen2022tolerance), we isolate employment protections and link them to labor supply, wage ranks, fertility, and opinion dynamics.
1.1 Related Literature
Research on sexual-orientation anti-discrimination laws (ENDAs) spans cross-sectional, causal state-level, and related policy literatures. Early cross-sectional work linked protections to narrower wage gaps but could not establish causality: klawitter1998effects use the 1990 Census and find no significant wage differences in protected places, while gates2009impact use the 2000 Census and report small premia. tilcsik2011pride found that resumes signaling LGB status received significantly fewer callbacks in localities without sexual orientation anti-discrimination laws. klawitter2011multilevel estimate multi-level earnings models with 2000 Census data, documenting systematic associations between state/local policies and earnings but not exploiting within-place policy changes over time. These past studies on anti-discrimination laws, while informative, fail to fully account for changes over time.
Quasi-experimental state-level studies strengthen identification. martell2013endas implement a state DiD design and find sizable wage convergence for gay men post-ENDAs, and Burn2018 adopt a triple-difference framework, still at the state level. A limitation of these state-only designs is that local ordinances, the likely first binding margin for many workers, are omitted; moreover, conventional TWFE estimators can be biased under staggered adoption and heterogeneous effects. To highlight these issues, Appendix B replicates a state-only specification and shows how excluding local ordinances can mask meaningful within-state variation, underscoring the value of incorporating local laws into the analysis.
Two adjacent strands inform mechanisms. First, policy shocks around marriage legalization affect household specialization and fertility: hansen2020labor show changes in labor supply following marriage legalization; hansen2022tolerance link tolerance to paid work among cohabiting gays and lesbians. These patterns match our household-reallocation results for women and the fertility margin we study. Second, mann2025antidiscrimination document mental-health benefits from anti-discrimination laws, suggesting channels beyond pecuniary outcomes and complementing our evidence on opinion dynamics.
Broader anti-discrimination evidence underscores that legal protections and enforcement can shift hiring and employment margins across protected classes: disability (audit evidence in ArmourButtonHollands2018; policy evidence in Button2018Expanding) and age (field-experiment evidence in NeumarkBurnButtonChehras2019). While these are not sexual-orientation ENDAs, they corroborate that statutory protections and their salience can meaningfully change labor-market opportunities.
@@ 2 Data
A common challenge in analyzing pay disparities between different types of same-sex and different-sex couples is the limited availability of high-quality data on sexual orientation, wages, and employment. We follow the literature in inferring sexual orientation by examining household relationship structure in the ACS (klawitter1998effects; gates2009impact; jepsen2017self). Specifically, we classify a person as being in a same-sex couple if they report a spouse or “unmarried partner” who is the same sex as themselves. [^4] The ACS measures self-reported “sex” rather than gender identity. Following the existing literature, we interpret this variable as a proxy for gender in couple-type inference, but this may misclassify some transgender or gender-nonconforming respondents. See, for example, gates2009impact and jepsen2017self. We exclude individuals whose sex or relationship status was allocated following gates2009same. To identify “unmarried partners,” the ACS asks each household member to report their relationship to the householder from a fixed list of categories. The category “unmarried partner” is defined by the Census Bureau as a person who has a close personal relationship with the householder. Importantly, the questionnaire provides a separate category for “housemate/roommate,” which reduces the likelihood that unrelated adults, such as different-sex roommates, are erroneously classified as unmarried partners.
@@
We use the 2005-2019 yearly ACS from ipumsusa. We use wages, defined as salaried wages from an employer. The ACS reports wage earnings in annual terms. We use the annual figures and convert them to hourly wage figures using variables on average weekly hours worked and weeks worked.555“Weeks worked” in the ACS is a categorical variable giving a range of weeks worked. We take the median value given in the range to compute weeks worked. Finally, we limit our sample to prime-age working adults and only examine those who are ages 25-65. We use 25 as the lower age cutoff to allow workers to complete college and have more fully developed human capital, which is typically thought of as an important factor for the differences in wages between LGB and straight workers (black2007economics).
@@
We obtained data on the passage of state anti-discrimination laws from LGBTMap.org, an LGBT advocacy group. The website gives information on which states passed anti-discrimination laws and when. We focus solely on sexual orientation anti-discrimination laws that give protection in employment. LGBTMap.org also provided incomplete data detailing the passage of local city and county laws with many cities missing years for the start of their anti-discrimination laws. We completed the dataset for the relevant years using old media reports, correspondence with local officials, and FOIA requests. Building on earlier city-level analyses such as klawitter2011multilevel, this dataset provides the first complete city-level dataset covering 15 years of sexual orientation anti-discrimination laws in the U.S. We focus on cities reported in the ACS and match those city laws with their corresponding counties to merge with the ACS. The ACS only reports county of residence for those in metro areas, so any rural counties are lumped together within a state.
We report the state, city, and county laws and the year they were enacted in Table 1.666Since August 2018, Pennsylvania has interpreted discrimination based on sexual orientation and gender identity as covered under the “sex” category of the Pennsylvania Human Relations Act, allowing LGBT individuals to file complaints for discrimination in employment, housing, education, and public accommodations. In Figure 2, we show which counties had sexual orientation anti-discrimination laws in 2005 and 2019, respectively. The enactment of anti-discrimination legislation is not random. The states that have these protections are generally considered friendlier to LGB workers than those without these laws and have a larger share of same-sex couples than those states without these laws. Also, many counties with anti-discrimination laws have large cities with a larger concentration of LGB workers than rural counties. However, this is not universal. For instance, Utah extended protection to LGB workers despite being a relatively conservative state, and certain liberal cities like Houston are noticeably absent from extending protection to LGB workers. [^7] Houston passed a sexual orientation and gender identity anti-discrimination ordinance, but it was only in effect for 3 months before being challenged. The law was put up to a public vote and lost, repealing the law. In Figure 2, we present the distribution of same-sex couples by state from 2005 to 2019, and in Table 2, we present the 10 counties with the largest share of same-sex couples. Unsurprisingly, the distribution of same-sex couples by state and by county is skewed toward more progressive states and counties with large cities in them that are known for having a large LGB population like San Francisco, the District of Columbia, New York City, and Boston.